A ball projected with velocity of10 m/s at angle of 30° with horizontal just clears two verticals poles each of height 1m find separation between the poles

Grade:12th pass

1 Answers

Shubham Pathy
5 years ago
We know equation of trajectory of a projectile
 
y=xtanm  - (gx^2/2(ucosm)^2).     Here m is angle of projection.
 
Substituting y=1and m=30° and g=10
We get 
(x^2)/15  -x/√3  + 1=0
 
=> X=2.29 ,6.52
 
So, distance from first hurdle = 2.29metre
Dr
Distance from second hurdle = 6.52 metre
 
So distance between the hurdles ,d=6.52-2.29=4.23metre